CAT Chat Minutes
February 9, 2004
General
Machine Stats - R. Klaffky:
Delivered beam: 245.24 hrs.
Percentage of scheduled time: 98.89
Downtime during period: 2.74 hrs.
Mean time between faults: 81.75 hrs.
Faults:
# 12 02/03 08:00 To 02/04 11:43 27.72 |
Power sag 89% [ComEd ] |
0.39 recovery, refill, resume top-up |
# 13 02/04 12:06 To 02/09 07:59 115.89 |
Int.dump: end of period |
0.00 |
Beamline Downtime: Period 1/29- 2/05/2004
Number |
Date |
Time |
Beamline |
Description |
Downtime |
Group |
Resolution |
000344 |
1/29/04 |
08:00 |
FM1 failure |
9days 9hrs |
VAC |
Replacement is underway (2/9) |
|
000329 |
1/29/04 |
08:30 |
Watchdog relay failure |
4:20 |
SI |
Re-wiring of relay by SI group |
|
000330 |
1/29/04 |
08:30 |
Closed beamline valve prevented station C from taking beam |
3:10 |
Beamline |
SI group assisted in correction of BLEPS issue |
|
000332 |
1/29/04 |
17:51 |
Beamline was taken offline to complete a portion of PSS validation |
0:38 |
SI |
Validation was completed |
|
000333 |
1/30/04 |
11:00 |
Watchdog relay failure |
0:30 |
SI |
Replacement of relay |
|
000334 |
1/31/04 |
20:33 |
LOVE controller set-point incorrect; slight water pressure fluctuation tripped high alarm set-point |
2:47 |
SI |
High alarm set-point was changed for c onsistency |
|
000336 |
1/31/04 |
12:00 |
Problem with drivers of L5-20 slits |
4:00 |
Beamline |
Beamline personnel corrected problem |
|
000337 |
2/2/04 |
13:15 |
CPU polarity could not be switched |
2:17 |
PS |
PS personnel replaced a bad communication cable |
|
000338 |
2/3/04 |
10:34 |
Door closed limit switch was not engaging properly; shutter could not be opened |
0:54 |
SI |
SI technician adjusted limit switch |
|
000339 |
2/3/04 |
22:35 |
BLEPS trips repeatedly closed shutter |
2:15 |
Beamline |
Beamline personnel corrected problem |
The following, from Greg Banks (AOD-Operations), explains status of 1ID:
The Vacuum group has everything leak tight, and is wrapping the mask for bake out. Mohan will need to go to ID-1 to verify that the proper mask is in place. The Vacuum group expects the work (bake- out, final leak-check) to be com pl eted tomorrow night between 08:00-10:00 p.m.
There will be repr esentatives from ASD Survey, Mec h an ical, and Water to button everything up when the vacuum work is complete. HP will verify shielding, the Ops will then close Zone A. Studies are planned to start at midnight Tuesday, and should start on time.
Operations Directorate: The issue of spare parts was discussed at today's meeting. We will ascertain that enough parts on-hand for the future.
ANL Domain Accounts: Automated email messages were sent informing of a password re-set requirement for the ANL domain account. You may change it using the ANL Credentials web page at https://credentials.anl.gov . Select the "Change Password Utility for Users" link. If you use Windows you may press CTRL+ALT+DEL and select the "Change Password" button.
Note that there are strict password security requirements that are enforced by the computer system. The rules are detailed on the web page where you change your password (see https://credentials.anl.gov ). If you have questions or have problems changing your password, please contact the division password administrator [Rachel Reed (x3656), Cindy Doran (x1783)] or the Account Services Office at 2-6969.
Laser Safety - E. Chang: Argonne Area Office (AAO) is currently reviewing our laser safety. We should expect to see them walking through the APS some time in late February. They will look at the various setups within APS.
ESAF Form - B. Glagola: Additional drop-down fields for the ESAF form are: GUP ID: (type in field); GUP, CAT, or PUP No. (choose from drop-down menu).